OpenStudy (anonymous):

I need help. Find the missing Values. Marked price Rate of Discount Discount Sale Price $800 _____ ___% $160 $________ (Simply your answer Type a whole number or decimal rounded to the nearest dollar as needed.)

OpenStudy (anonymous):

It's very simple, really. What's 10% of $800? (:

OpenStudy (anonymous):

0.010* 800= 8

OpenStudy (anonymous):

Jon_persson do i have too multiply 0.10* 800

OpenStudy (anonymous):

That's 1% of 8, but what's 10%? Yes, multiply 0.10 by 800.

OpenStudy (anonymous):

it = 80

OpenStudy (anonymous):

so 80 is therate of discount

OpenStudy (anonymous):

How would you find the sales price

OpenStudy (anonymous):

No, $80 is not the rate of discount. It's what the discount in dollars would be if the rate would have been 10%.

OpenStudy (anonymous):

ok. so How would I figure out the rate of discount

OpenStudy (anonymous):

Well, if $80 USD is 10% of the price, what is $160?

OpenStudy (anonymous):

the sales price isnt 80 dolalrs,

OpenStudy (anonymous):

im confuse.

OpenStudy (anonymous):

\[800\div160=5\] \[100\div5=20\] \[800-160=640\]

OpenStudy (anonymous):

so the sales price is 640 the rate discount is 20 %

OpenStudy (anonymous):

So, you're trying to find how many percent of 800 dollars $160 is. To do this, one way to go about it is to figure out what 10% of 800 dollars is. \[0.10*800=80\] So 10% equals 80 dollars. What would 20% be, in dollars? \[0.20*800=160\] 160 dollars. Therefore, the rate of discount is 20%. Now to figure out the sales price. \[800-160=640\] Another, perhaps more scientific/mathematical way to go about it is to simply divide 800 by 160. \[800\div160=5\] So 160 is one fifth of 800. Now, to find out the percentage of one fifth, simply divide 100 (100% of something) by 5 (one fifth of that 100%) \[100\div5=20\] Therefore, one fifth of something is the same as 20% of something. So, the rate of discount is 20%. Now that that is done, all we have to do is subtract. \[800-160=640\]

OpenStudy (anonymous):

nabizag is correct.